There is a solid ball that weighs 1000g. The radius of the ball is 5 cm.

1. What is the density of the ball? (ππ=3)    in kg/m3    (no unit needed in the answer )  

2. If you cut a smaller ball, which has a radius of 1 cm, from this big ball. What is the mass left? (\piπ=3)    in kg    (no unit needed in the answer )  

3. What is the volume ratio of those two balls? (small : big)  
Answer #1

1) We know that density d=M/V ....1

here M=1000g =1.0 kg , V=4\prodR3/3 R=5 cm =0.05 m ,\prod=3 , so V=4\prod(0.053)/3 =5*10-4 m3

d=1/5*10-4

d=2000 kg/m3

2) volume of the small ball (R=1cm=0.01 m) Vs=4\prod(0.013)/3 =*10-6 m3

mass of the ball m=dV =2000*4*10-6

m=8000 *10-6

m=8*10-3 kg

left mass =1-0.008=0.992 kg answer

3)

Vs/V=4.*10-6/5*10-4

=0.8*10-2

=0.008 answer